# Advancing Multi-Organ Disease Care: A Hierarchical Multi-Agent Reinforcement Learning Framework

### v1

@@ Introduction
RQ3: How does the learned policy compare to the clinician policy in the context of multi-organ treatment decisions?
We assess the alignment between clinician and learned policy using correlations among all treatments within each policy. For granular comparisons, these correlations were further categorized by low, medium and high SOFA severity levels at the time of intervention (see Figure 4). While exact concordance is not expected since survival trajectories could often include non-optimal actions, proposed policy generally aligns with clinician decisions linked to patient survival and deviates from those associated with mortality.
@@
RQ4: How effective are the proposed dual-layer state representations and cross-agent communication mechanisms?
@@
We measured the contributions of two core components of our proposed solution: dual-layer state representations and cross-agent communications. We trained three variations of our hierarchical model: one without inter-agent communications (Prop-NoC), one without the proposed state representations, using raw feature vectors (Prop-NoSR), and without both (Prop-NoC-NoSR). All models showed degraded performance in mortality rates and CWPDIS values (see Tables 1 and 2). Compared to our full proposed model, we see an increase of 19.6%, 27.2%, 33.4% in mean mortality for Prop-NoC, Prop-NoSR, and Prop-NoC-NoSR, respectively. This indicates the critical role of these components in handling the task complexity.
@@
Table 2: Off-policy evaluation metrics of the proposed model, without inter-agent communication and state representations.
Model ESS V Mortality (%)
Prop-NoC 4775 26.27 10.54 ± 0.27
Prop-NoSR 1020 26.13 11.21 ± 0.33
Prop-NoC-NoSR 3633 24.80 11.75 ± 0.50
Conclusions
@@
We introduce a hierarchical multi-agent reinforcement learning framework for the complex and first-of-its-kind multi-organ treatment recommendations, setting a new benchmark in clinical decision support systems. Mimicking real world collaborative clinical settings, our solution effectively decomposes the complex treatment process into a clinically meaningful hierarchy of subtasks. Each subtask is managed by specialized agents operating within dedicated subspaces. It supports both independent and cooperative agent functionality through robust inter-and intra- organ communications. Moreover, a dual-layer state representation technique is proposed to support advanced contextualization needed at multiple levels in the hierarchy. We evaluate our solution on the non-standardized and multi-dimensional sepsis treatment recommendation. Comprehensive quantitative and qualitative evaluation showed that our solution consistently outperformed baselines, significantly improving the patient survival and effectively managing task complexity. Furthermore, learned policy closely followed successful clinical treatment patterns, deviating only when beneficial, thus enhancing the reliability of the policy. The inherent flexibility and scalability of our solution allow it to be expanded to a broader range of treatments and organ systems, and even for complex decision-making scenarios beyond healthcare.
Technical Appendix
Methodology
RL Actions:
@@
For each 4-hour window, we computed the total dosage for each treatment by multiplying its infusion rate with the overlapping duration, and adding any IV push volumes. Treatments with multiple drugs were converted into their standard equivalents (i.e., vasopressors into norepinepherine (Kotani et al. 2023), S2 to fentanyl (McPherson et al. 2010), and diuretics to furosemide (Konerman et al. 2022)) (see Table 1).
@@
Figure 5: Baseline Performances omitted due to space limitations from main text. Dosage differences (x-axis) versus mortality (y-axis) for individual agents in the hierarchy, with colors blue, red, and green denoting different organ systems.
@@
Table 1: Summary of treatments, clinical components, purposes, and number of samples per treatment
@@
Vasopressors Norepinephrine, Epinephrine, Dopamine, Vasopressin, Phenylephrine Increase blood pressure for improved organ perfusion; stabilize hemodynamics 55,591
Renal Diuretics Furosemide, Bumetanide Reduce blood pressure and edema; maintain fluid balance 17,279
Dialysis Active renal replacement therapy Waste product removal, maintain electrolyte and fluid balance; support kidney function 7,135
RL Reward:
In addition to the mortality based terminal reward ($+R/-R$), in line with the literature, we use a clinically guided intermittent reward function $R_{im}(s_{t},s_{t+1})$, calculated from the transition from $s_{t}$ to $s_{t+1}$ based on immediate impact on person’s health post-treatment (Raghu et al. 2017).
Table 2: List of model features
@@
Demographics (9) Age, Elix., Shock index, SOFA, GCS, Weight, SIRS, Gender, Readmission
@@
Lab values (24) Albumin, pH, Calcium, Glucose, Hb, Magnesium, WBC, Creatinine, Bicarbonate, Sodium, CO2, Lactate, Chloride, Platelets, Potassium, PTT, PT, AST, ALT, BUN, INR, Ionised calcium, Total bilirubin, Base excess
@@
  • Abbreviations- INR: International Normalized Ratio; PT: Prothrombin Time; PTT: Partial Thromboplastin Time; SIRS: Systemic Inflammatory Response Syndrome; ICU: Intensive care unit; WBC: White blood cell; Temp.: Temperature; GCS: Glasgow Coma Scale; Resp.: Respiratory rate; HR: Heart rate; SBP: Systolic blood pressure; MBP: Mean blood pressure; DBP: Diastolic blood pressure; Hb: Hemoglobin; Elix.: Elixhauser score.
Experiments
Dataset:
We used 48 physiological features including demographics, lab values, vital signs and intake/output events (see Table 2) and aggregated data into 4-hourly windows using mean or sum as appropriate. Missing values were imputed with the last available data from previous windows. Binary features were normalized to -0.5 and 0.5, and normally and log-normally distributed features to 0-1 range. The study included 25,492 survivors (41.7% female) and 4,948 deceased patients (44.0% female).
Reproducibility and Model Parameters:
Source codes along with tuned parameters and architectures are uploaded as technical appendix with the submission and available online as an anonymous repository [^3] https://anonymous.4open.science/r/HMARL-PyTorch-022B/. All models are tested on the same holdout set. We experimentally set $\gamma$ to 0.99 and $k$ to 8. The sepsis cohort can be replicated using the provided SQL scripts and Python codes.
Mortality vs. Difference in Recommended Dosage:
@@
Individual Agent Performance:
Figure 6 shows mortality verses dosage differences plots of individual agents.
Figure 6: Performances of individual agents in the hierarchy. Dosage differences (x-axis) versus mortality (y-axis) for individual agents in the hierarchy, with colors blue, red, and green denoting different organ systems.
Ethical Considerations:
Similar to RL based Clinical Decision Support Systems (CDSSs) proposed in the literature for various diseases including sepsis management (Komorowski et al. 2018; Raghu et al. 2017; Saria 2018), our solution is a human in the loop CDSS; where clinicians could utilize the data driven decisions provided by the intelligent agents as auxiliary inputs before making the final recommendations. In addition to the final action recommendations from the CDSS, the $Q(s,a)$ values available for all possible actions ($a$) could be used to quantify and compare the quality of the available treatment options for a given patient, with respect to his long term survival. The proposed solution allows clinicians to collaborate with the intelligent agent that combines the experiences of successful clinical decisions in the past, but leaves the full control of the final decision to the clinician. Furthermore, in line with the literature, all models were trained and evaluated using publicly available retrospective data, and no clinical trials were conducted. Therefore, the proposed solution does not raise ethical concerns.

### v2

@@ 3.3.2. RQ2: Do individual agents learn effective local policies?
3.3.3. RQ3: How does the learned policy compare to the clinician policy in the multi-organ treatment decision context?
We further assess how our learned policy aligns with clinician behavior in multi-organ intervention contexts. Specifically, we compare the pairwise treatment correlations across six intervention types between the clinician and the proposed policy. These comparisons are stratified by SOFA-based patient severity (i.e., low, medium and high) and survival outcome (i.e., survived vs. deceased) (see Figure 5). Note that exact alignment with clinician practice is neither expected nor optimal since it is well documented that clinicians often make suboptimal decisions under uncertainty in complex disease management. However, alignment with positive outcomes (i.e. survival) and divergence in high-risk cases (i.e., non-survivors) would showcase a robust and adaptive policy. It is evident that our policy closely replicated the positive treatment synergies observed in survivors across different severity levels, while showcasing reduced correlation with deceased cohorts. This indicates that the proposed model was able to recognize effective clinical patterns and also learned to avoid suboptimal treatment combinations associated with poorer outcomes. These findings showcase the policy’s ability for adaptive decision-making based on patient state severity that could lead to superior outcomes.
Figure 5. Correlation matrices of 6 treatment types within clinician and proposed policies across low, medium, and high SOFA severity levels, categorized by (a) survived and (b) deceased outcomes.
@@
3.3.4. RQ4: How effective are the proposed dual-layer state representations and cross-agent communication mechanisms?
@@
We measured the contributions of two core components of our solution: dual-layer state representations and cross-agent communications. We trained three variations of our hierarchical model: without inter-agent communications (Prop-NoC), without the proposed state representations, using raw feature vectors (Prop-NoSR), and without (Prop-NoC-NoSR). All models showed degradation in mortality rates and CWPDIS values (see Tables 2 and 3). Compared to our full model, we see an increase of 27.2%, 19.7%, 33.4% in mean mortality for Prop-NoC, Prop-NoSR, and Prop-NoC-NoSR. These results demonstrate the critical role of these components.
@@
Table 3. Off-policy evaluation metrics of the proposed model, without inter-agent communication and state representations.
Model V Mortality (%)
Prop-NoC 18.54 12.65 $\pm$ 0.27
Prop-NoSR 19.01 11.92 $\pm$ 0.36
Prop-NoC-NoSR 16.18 13.27 $\pm$ 0.31
4. Conclusions
@@
We introduce a hierarchical multi-agent reinforcement learning framework for the complex and first-of-its-kind multi-organ treatment recommendations, setting a new benchmark in clinical decision support systems. Mimicking real world collaborative clinical settings, our solution effectively decomposes the complex treatment process into a clinically meaningful hierarchy of subtasks. Each subtask is managed by specialized agents operating within dedicated subspaces. It supports independent and cooperative agent functionality through robust inter-and intra- organ communications. Moreover, a dual-layer state representation technique is proposed to support advanced contextualization needed at multiple levels in the hierarchy. We evaluate our solution on the non-standardized and multi-dimensional sepsis treatment recommendation problem. Comprehensive quantitative and qualitative evaluation on internal and external validation datasets showed that our solution consistently outperformed baselines, significantly improving the patient survival and effectively managing task complexity. Furthermore, learned policy closely followed successful clinical treatment patterns, deviating only when beneficial, thus enhancing the reliability of the policy. The inherent flexibility and scalability of our solution allow it to be expanded to a broader range of treatments and organ, and also for complex scenarios beyond healthcare.
Appendix A Methodology
A.1. RL Actions:
@@
For each 4-hour window, we computed the total dosage for each treatment by multiplying its infusion rate with the overlapping duration, and adding any IV push volumes. Treatments with multiple drugs were converted into their standard equivalents (i.e., vasopressors into norepinepherine (Kotani et al., 2023), S2 to fentanyl (McPherson et al., 2010), and diuretics to furosemide (Konerman et al., 2022)) (see Table 4). Thresholds used for action discretization can be found in Table 5.
A.2. RL States:
State variables were chosen based on clinical relevance and concurrent availability in the MIMIC-IV and AmsterdamUMC datasets (see Table 6). The data was aggregated into 4-hourly windows using mean or sum as appropriate. Missing values were imputed using the ’last-observation carried forward’ method (LOCF). Binary features were normalized to -0.5 and 0.5, and normally and log-normally distributed features to 0-1 range.
@@ A.2. RL States:
Figure 6. Baseline performances omitted due to space limitations from main text. Dosage differences (x-axis) versus mortality (y-axis) for individual agents in the hierarchy, with colors blue, red, and green denoting different organ systems.
@@
Table 4. Summary of treatments, clinical components, purposes, and number of samples per treatment
@@
Vasopressors Norepinephrine, Epinephrine, Dopamine, Vasopressin, Phenylephrine Increase blood pressure; stabilize hemodynamics 55,591
Renal Diuretics Furosemide, Bumetanide Reduce BP and edema; maintain fluid balance 17,279
Dialysis Active renal replacement therapy Waste removal, support kidney function 7,135
Table 5. Action level thresholds (4-hourly doses)
Treatment 0 1 2 3 4
IV Fluids (ml) 0 0.00–56.17 56.17–227.50 227.50–530.93 >>530.93
Vasopressor (µg/kg) 0 0.00–9.40 9.40–20.66 20.66–44.42 >>44.42
S1 (mg/kg of propofol) 0 0.00–3.40 3.40–6.01 6.01–9.53 >>9.53
S2 (µg/kg of fentanyl) 0 0.00–0.65 0.65–2.08 2.08–4.30 >>4.30
Diuretics (mg of furosemide) 0 0.00–20.00 20.00–160.00 160.00–902.10 >>902.10
Dialysis Off On
A.3. Training Process
We describe the overall training process in Algorithm 1.
Algorithm 1 Generalized Training Process for HMARL Treatment Dynamic Treatment Recommender Solution
0:
0:$W$: a set of organ systems {i.e., {Neu, Car, Ren}}
0:$S$: learned state representations, $S=\{s_{t}^{Rt},s_{t}^{o}\mid\forall o\in W\}$
[…]
0:$r$: reward signal
[…]
1:Initialize: $Q(s,a)$ for:
2:     Root agent: $M_{Rt}$
3:     Single-organ agents: $M_{o}$ for all $o\in W$
4:     Treatment-level agents: $M_{o}^{treat}$ for all $o\in W,\forall treat\in A_{o}$
5:     Mixed-treatment agents: $M_{o}^{Mix}$ for all $o\in W$
6:     Mixed-organ agents: $M_{OMix}$
7:Step 1: Train the Root Agent $M_{Rt}$
8: Input: $s_{t}^{Rt}$ {Unified state representations}
9: Output: $a\in A_{Rt}$ where $A_{Rt}=\{A_{Rt}^{0},A_{Rt}^{OMix}\}\cup\bigcup_{o\in W}A_{Rt}^{o}$
10:Step 2: Train Single-Organ System Agents
11:for $o\in W$ do
12:  Train $M_{o}(s_{t}^{o},a)$ on $a_{c}\in A_{o}$
13:  Output: $a_{o}^{master}\in\bigcup_{i=1}^{|A_{o}|}a_{o}^{master_{i}}$
14:  for $treat\in A_{o}\setminus A_{o}^{Mix}$ do
15:   Train $M_{o}^{treat}(s_{t}^{o},a)$ on $a_{c}\in A_{o}\setminus A_{o}^{Mix}$
16:   Output: $a\in\bigcup_{d\in D}a_{o}^{treat,d}$
17:  end for
18:  Train $M_{o}^{Mix}(s,a)$ on $a_{c}\in A_{o}^{Mix}$, where:
19:  $s=\text{concat}(s_{t}^{o},a_{o}^{treat}\mid\forall treat\in A_{o}\setminus A_{o}^{Mix})$
20:  Output: $a\in\bigcup_{(treat,d)\in|A_{o}|\times D}a_{o}^{treat,d}$
21:end for
22:Step 3: Retrieve Single-Organ System Actions
23:for $o\in W$ do
24:  Select $a_{o}^{master}$ from $M_{o}$
25:  Retrieve $G_{o}=\{a_{o}^{treat_{1}},\dots,a_{o}^{treat_{n}}\}$, where $n=|A_{o}|$
26:end for
27:Step 4: Train the Mixed-Organ Agents using QMIX
28: Agents: $M_{OMix}(s,a)$ and $M_{OMix}^{o}$ for all $o\in W$
29: Input: $s=\big{[}s_{t}^{y},G_{y}\mid\forall y\in W\big{]}$
30: Output: Mixture of treatments across organ systems: $A_{t}^{OMix}$
31:Step 5: Phase-2 Training
32: See Equation (1) for detailed updates
Table 6. List of model features
@@ A.3. Training Process
Demographics (8) Age, Shock index, SOFA, GCS, Weight, SIRS, Gender, Readmission
@@
Lab values (21) Albumin, pH, Calcium, Glucose, Hb, Magnesium, WBC, Creatinine, Bicarbonate, Sodium, Lactate, Chloride, Platelets, Potassium, PTT, AST, ALT, BUN, Ionised calcium, Total bilirubin, Base excess
@@
  • Abbreviations- PTT: Partial Thromboplastin Time; SIRS: Systemic Inflammatory Response Syndrome; ICU: Intensive care unit; WBC: White blood cell; Temp.: Temperature; GCS: Glasgow Coma Scale; Resp.: Respiratory rate; HR: Heart rate; SBP: Systolic blood pressure; MBP: Mean blood pressure; DBP: Diastolic blood pressure; Hb: Hemoglobin
Appendix B Experiments
B.1. Mortality vs. Difference in Recommended Dosage:
@@ Appendix B Experiments
B.2. Ethical Considerations:
Similar to RL based Clinical Decision Support Systems (CDSSs) proposed in the literature for various diseases including sepsis management (Komorowski et al., 2018; Raghu et al., 2017; Saria, 2018), our solution is a human in the loop CDSS; where clinicians could utilize the data driven decisions provided by the intelligent agents as auxiliary inputs before making the final recommendations. In addition to the final action recommendations from the CDSS, the $Q(s,a)$ values available for all possible actions ($a$) could be used to quantify and compare the quality of the available treatment options for a given patient, with respect to his long term survival. The proposed solution allows clinicians to collaborate with the intelligent agent that combines the experiences of successful clinical decisions in the past, but leaves the full control of the final decision to the clinician. Furthermore, in line with the literature, all models were trained and evaluated using publicly available retrospective data, and no clinical trials were conducted. Therefore, the proposed solution does not raise ethical concerns.
B.3. Reproducibility and Model Parameters:
Source codes along with tuned parameters and architectures are uploaded with the submission and available online as an anonymous repository [^1] https://anonymous.4open.science/r/HMARL-PyTorch-022B/. All models are tested on the same holdout set. We experimentally set $\gamma$ to 0.99 and $k$ to 8. Data for the sepsis cohort can be replicated using the provided SQL scripts and Python codes.